tourists enter a popular game reserve at an average rate of one every five minutes. what is the probability that it takes more than twenty minutes until the first tourist arrives?


1
Expert's answer
2021-06-08T12:49:18-0400

Let "X=" the number of tourist arrives: "X\sim Po(\lambda t)"

Given "\lambda=1, t=4"


"P(X=0)=\dfrac{e^{-\lambda t}\cdot(\lambda t)^0}{0!}=e^{-1(4)}\approx0.018316"

The probability that it takes more than twenty minutes until the first tourist arrives is

"0.018316."